OpenStudy (anonymous):

find the inverse laplace transform of (3/(2s+5)^3).

OpenStudy (amistre64):

hmm factor out the 2 is my first thought; try to get this into: 1/(s+k)^3

OpenStudy (anonymous):

I'm confused on how to do this problem in general. i don't know what Laplace transform is

OpenStudy (amistre64):

if you dont know what Laplace transform is, then you reallya re asking the wrong question to begin with .....

OpenStudy (amistre64):

the Laplace transform is an integral transform that .. well, transforms a function from an integration to an algebra problem

OpenStudy (amistre64):

there is a table of Laplace transforms that we try to get this thing to look like in order to undo it back into what it looked like to begin with

OpenStudy (amistre64):

my idea for the inverse is to basically format it into something invertable

OpenStudy (amistre64):

\[\frac{3}{(2(s+\frac{5}{2}))^3}\] \[\frac{3}{8(s+\frac{5}{2})^3}\] \[\frac{2*\frac{3}{2}}{8(s+\frac{5}{2})^3}\] \[\frac{3}{16}L\left\{\frac{2}{(s+\frac{5}{2})^3}\right\}\]

OpenStudy (amistre64):

i got no idea how right t is tho; so id have to chk my resulte wolfs with th

OpenStudy (amistre64):

laptop mistype me up lol

OpenStudy (anonymous):

why do the initial variable x become t?

OpenStudy (anonymous):

is it supposed to show that it's not the original problem?

OpenStudy (amistre64):

the Laplace transform takes a function of one variable; namely t and transforms it into a function of another variable, namely s ; but i chose x in this place cause the wolf was being nitpicky and wanted to read the s as seconds instead of a variable ....

OpenStudy (anonymous):

so after you formed the function to a more workable problem, you just plug into a formula?

OpenStudy (amistre64):

the laplace tables tell me: L{t^n} = n!/s^(n+1) so if I can get the function you provided into this format, I can undo (or invert) the Laplace transform to get back to the original function i just wasnt to sure what i should do to the extra +5/2 in the bottom

OpenStudy (amistre64):

the site seems buggy tonight ...

OpenStudy (anonymous):

what do you mean by the original function?

OpenStudy (anonymous):

and how do you get from L (2/(s+5/2)^3) to plug into the form n!/s^(n+1)

OpenStudy (anonymous):

where is n?

OpenStudy (amistre64):

since the underside has ^3; which matches up to ^(n+1) ; n has to equal 2

OpenStudy (amistre64):

2*3/2 = 3 so I just rewrote the top to my advantage

OpenStudy (amistre64):

same with the bottom; i factored out a 2 down there

OpenStudy (amistre64):

pull the constants out that aint needed and your set to undo it

OpenStudy (anonymous):

so can we completely ignore constants?

OpenStudy (amistre64):

the Laplace is a linear transform, meaning it has the following properties L{a+b} = L{a} + L{b} and for any constant k, L{ka} = k L{a} so the constants can be pulled out as concerns the actual operations

OpenStudy (anonymous):

oh that makes sense!

OpenStudy (amistre64):

as such, we can format these things by using constants to get them to a proper form that can be inverted

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

youre welcome